1/65
Looks like no tags are added yet.
Probability
a number between 0 →1 that quantifies the likelihood of an event occurring
Frequentist
the fraction of times something happens if you repeat the situation forever
Bayesian
the degree of belief, a number measuring how confident you are between 0 → 1
Probability calculation clasical
P(a)= (# of elements in A) / (# of elements in S)
Relative frequency (empirical method)
when outcomes are not equily likely. we estimate the probability from the data.
Probability calculation empirical method
P(a)= (#of times A occurred)/(# of trials)
Probability calculation empirical method (excel)
=COUNTIF(A2:A601, "Late")/COUNTA(A2:A601)
Subjective method
When there is no repeatable experiment and little or no data, a probability is assigned by reasoned judgment
Assigning Probabilities requirements
Every probability is between 0 and 1
The probabilities of all the outcomes in the sample space add up to 1
Mutually Exclusive
two events cannot happen at the same time
Union
“or” the probability of hears or Spades
requires addtion
Intersection
“and” the probability of hearts and spades
requires multipulcation
Complements P(Ac)
The remainder, everything thats not the specific scenario
Complements calculation
P(Ac)=1-P(A)
De Morgan’s Law
the complement of a union is intersection and the complement of an intersection is union
Addition Rule (mutually exclusive)
P(AuB)= P(a)+P(B)
Addition rule (not mutualy exclusive)
P(AuB)= P(a)+P(b)-P(A⋂B)
Multiplication rule (dependent events)
P(A⋂B)= P(a|B) *P(b)
Multiplication rule (indepent events)
P(A⋂B)= P(a) * P(b)
Independent events
two events are independent if the probability of occurrence by one event is not effected by the occurrence of the other event
Conditional probability
measures the likelihood of an event occurring given that another event has already happened.
Conditional probability calculation
P(A|B)= (P(A⋂B))/(P(B))
Law of total probability
P(a)= P(A⋂B)+ P(A⋂Bc)
or
P(a)= P(A|B)P(B)+P(A|Bc)P(Bc)
contingency table

Bayes Rule
P(A|B)= (P(B|A)P(A))/(P(B)
counting rules (calculation)
p= (# elements in event)/(# elements in S)
Factorial
counts how many ways a group of objects can be organized
Fundamental counting principle
Ex: a restaurant offers 3 appetizers, 4 entrees, and 2 desserts. The number of different three course meals is 3×4×2=24
Factorial calculation
N! → x*(x-1)*(x-2)…1
combinations
counts the number of ways to choose x objects from a total of n objects
combinations formula (no repition)
c(n,x)= n!/((n-x)!x!)
combinations formula (with repition)
C(n+x-1,x)= (x+n-1)/(n-1)!x!
Permutation
also counts the number of ways to choose x objects from a total of n objects. However, the order in which the objects are listed does matter.
permutation calculation (no repetion)
P(n,x)= n!/(n-x)!
Permutation calculation (with repetition)
p(n,x)=nx
Contingency table in excel
Pivot table, place one variable in the rows area, the other in the columns area, and any field in the values area as a count
Factorial in excel
=FACT(number)
Combinations in excel
=Combin(number, number_chosen) (no repetition)\
=Combina(number, number_chosen) (no repetition) with repition
Permutation in excel
=Permut(number,number_chosen) no repetition
=permutationa(number,number_chosen) with repetition
Expected Value
a measure of central location
Expected value calculation
E(x)= sum xf(x)
Variance
summarizes the deviation of the values of the random variable from the mean.
variance calculation
var(x)= E[(x-E(x))2)]=E(x2)-[E(x)]2
E(x2)- m²
uniform
is a probability distribution that assigns equal probability to each outcome
uniform calculation
f(x)=1/n
e(x)= (sum x)/n
var(x)= sum(x-m)²/n
Binomial distribution
counts the true # of successes in n trials of an experiment
x→ captures the # of successes
n→ number of trials
every outcome can be classified as success or failure
the probability of success is P and fixed
trials are independent
binomial distribution calculation
f(x)= Cxn (px)(1-p)n-x
E(x)=np
var(x)=np(1-p)
hypergeometric distribution
a probability distribution that describes the outcome of drawing a sample from a population without replacement.
trials are not indecent
probability of success changes from trial to trial
hypergeometric calculation

Poisson Distribution
estimates the number of successes (x) over a specified interval of time or space
An experiment satisfies a Poisson process if:
The number of successes with a specified time or space interval equals any integer between zero and infinity.
The number of successes counted in non-overlapping intervals are independent.
The probability of success in any interval is the same for all intervals of equal size and is proportional to the size of the interval.
poisson calculation
F(x)= (uxe-u)/x!
u- arival rate (average successes in the given time)
Binomial in excel
=binom.dist(x,n,p,FALSE)
poisson in excel
=Poisson.dist(x,mean, FALSE)
hypergeometric in excel
=hypgeom.dist(x,n,r,N,FALSE)
=HYPGEOM.DIST(sample_s, number_sample, population_s, number_pop, cumulative)
Probability mass function (PMF) in excel
set the last cumulative argument to FALSE
Cumulative probabilities in excel
set the last argument to TRUE
continuous random variables
a variable that can take on an uncountably infinite number of possible values within a given range or interval
uniform distribution
The uniform distribution assigns equal density across an interval [a,b]
the uniform distribution (calculation)
f(x)= 1/(b-a) (a≤x≤b)
e(x)= (a+b)/2
var(x)= (b-a)2/12
where a is the lower limit ans b us the upper limit f(x) 0 outside [a,b]
normal distribution
It is symmetrical about the mean u.
The mean is at the middle and divides the area of the distribution into halves.
The total area under the curve is equal to 1.
The distribution is completely determined by its mean and standard deviation.
The standard normal has a mean of 0 and a standard deviation of 1. Otherwise, it has the exact same properties as the normal distribution.
exponential distribution
is useful in computing probabilities for the time it takes to complete a task. It describes the time between events in a Poisson process
Exponential distribution calculation
f(x)= (1/u)e-x/u
Triangle distribution
is characterized by a single mode (the peak of the distribution) and two boundaries. It is often used in situations where the lower and upper bounds of a potential outcome are known, but the exact likelihood of the outcome is uncertain.
Triangular distribution formula

Normal distribution (excel)
Norm.dist(x,mean,standard_dev,cumulative)
set cumulative to TRUE for P(X≤x)
set cumulative to FALSE for the density
for standard normal (u=0,ō=1) use Norm.s.dist(z, cumulative)
to find value given a probability use Norm.inv(probability,mean,standard_dev) or Norm.S.INV(Probability)
Exponential excel
=expon.dist(x,1/u, cumulative)